OpenStudy (anonymous):

Simplify: sqrt c (3 + 2sqrt 5) 5 sqrt 5c C sqrt 3 + c sqrt 10 3 sqrt c + 2 sqrt 5c 3 sqrt c + 2c sqrt 5

OpenStudy (ajprincess):

Use distributive property a(b+c)+ab+ac Here a=sqrtc b=3 c=2sqrt5 Does that help? @Solmyr

OpenStudy (anonymous):

srry had to do something

OpenStudy (anonymous):

Sqrt c(3 + 2 sqrt 5) + sqrt c(3) + sqrt c(2 sqrt 5)

OpenStudy (phi):

it is best to put the numbers in front of the sq rt sign, it's less confusing Sqrt c(3 + 2 sqrt 5) = 3sqrt c + 2 sqrt c sqrt 5 you should know that \[ \sqrt{a} \cdot \sqrt{b} = \sqrt{a\cdot b} \] That rule means you can combine sqrt(c)*sqrt(5)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i was thinking that the answer was A and still thinking because of what u showed

OpenStudy (anonymous):

ok

OpenStudy (phi):

I thought A was B. But A is not correct. the answer is 3sqrt c + 2 sqrt c sqrt 5 or a simplified version of this

OpenStudy (anonymous):

ok so D??

OpenStudy (phi):

rewrite \[ 3\sqrt c + 2 \sqrt c \sqrt 5\] by putting the c and 5 inside the same sqrt

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you have time to help me with two more?

OpenStudy (phi):

first, did you finish this one? put the c and 5 inside the same sqrt \[3\sqrt c + 2 \sqrt c \sqrt 5\]

OpenStudy (phi):

you can do that only for the 2nd term.

OpenStudy (anonymous):

3 sqrt c + 2 sqrt 5c

OpenStudy (phi):

ok, that should match one of your choices

OpenStudy (anonymous):

C

OpenStudy (phi):

so to do this, you distributed and then you combined two sq rt's

OpenStudy (phi):

please post your other question separately

OpenStudy (anonymous):

ok